Show how to get the reducible and irreducible representation for the vibrations of PF5 (D3h config).

Then show all the IR/Ramen active models

((Inorganic Chem))
Answer #1

pls chek

e 20 3c₂ ob 25 3oy 2 4 1 3 4 unshitted 6 aton 0 - 2 4 - 2 4 Redumble 18 Representation. Da b & 263 (2) 18 lo hlag) 253 1 + 1DA= 1/[ 18 tot 6 - 4 +X-x] [ 18 -18720 A2 = 11/12 [is to to – 4 +4 +12] DE = 136+0 -0 -8 – 4 to] =/2 [24] = 2 E 2AT A +